OpenStudy (anonymous):

ln e^7 Simplify without a calculator

OpenStudy (anonymous):

hi again apparently the answer is 7 but i'd like an explanation as to why if you can help me out.

OpenStudy (jdoe0001):

well let's see ok

OpenStudy (aum):

\[ \ln(a)^b = b * \ln(a) \\ \ln(e)^7 = 7 * \ln(e) \\ \ln(e) = 1 \\ \]

OpenStudy (jdoe0001):

\(\large \bf ln(e^7)\implies log_{\color{red}{ e}}{\color{blue}{ e}}^7=\square \implies {\color{red}{ e}}^\square ={\color{blue}{ e}}^7\)

OpenStudy (jdoe0001):

see what \(\square\) equals to?

OpenStudy (anonymous):

7 im guessing

OpenStudy (jdoe0001):

well, if the base is the same... and the EQUATion is true... then the exponents need to equal each other :)

OpenStudy (anonymous):

wait the base is the red e or the blue e?

OpenStudy (jdoe0001):

hmmm the base is the red one

OpenStudy (anonymous):

okay that makes sense also ln is that just logarithm? If so on the calculator which button would i use if it has both "Log"and "Ln"

OpenStudy (jdoe0001):

yeap is the "natural logarithm" or \(\large ln\iff log_e\)

OpenStudy (jdoe0001):

keep in mind the "log cancellation rule" thus when the log base and the "base of the number inside" match, then \(\Large \bf log_{\color{brown}{ a}}{\color{brown}{ a}}^x=x\)

OpenStudy (anonymous):

Ohhh wow! Cool! Thanks so much!

OpenStudy (jdoe0001):

yw